how do you factor this polynomial?
y^3-y^2+y-1
the answer in the back of the book says (y-1) (y^2+1) but i don't know how they got that.

Answer by shani(1)

that answer is correct. simply foil. first, outer, inner, last.
you get (y)(y^2)= y^3
then you get (y)(1)=y
then (-1)(y^2)= -y^2
and lastly (-1)(1)= -1
put them in decending power and you come up with y^3-2y^2+y-1